# Risk-Neutral Pricing Model of Uniswap Liquidity Providing Position: A Stopping Time Approach

### v2

@@ IV Sensitivity Analysis
Our subsequent Greeks analysis will be conducted based on two sets of market parameters corresponding to different market conditions. One scenario features high transaction fees and high market volatility, referred to as ’good market situation’, as these conditions allow Uniswap V3 positions to accrue more fees. The other scenario involves low transaction fees and low market volatility, termed ’good market situation’.
Taking the European LP option pricing model as an example, we present the valuation discrepancies between pricing under good/bad market conditions and the payoff model at different price levels. The good/bad market conditions referenced here are based on the market parameters used in Table I and Table I. The range order price interval is set at (0.8,1.2). As shown in Figure 3, under the calculations based on the European LP option pricing model, there is a significant difference between the current Uniswap V3 payoff value and its actual present value. This is because the current value of a Uniswap V3 position is not limited to the extractable payoff; it also incorporates the discounted value of future fees that can still be collected if the position is not immediately withdrawn.
In scenarios with relatively low volatility and low transaction fees, the analysis shows that if the Uniswap V3 position is not withdrawn now, its future value will depreciate, and the discounted value of future fees will be insufficient to offset the loss caused by the position’s depreciation, as depicted by the green line in the Figure 3. Conversely, under good market conditions with higher volatility and transaction fees, the orange line in the Figure 3 demonstrates that the actual value of the Uniswap V3 position exceeds its payoff value.
@@
IV-A Delta Analysis
In historical Uniswap V3 trading practices, to mitigate the uncertainty of future returns, delta hedging is often employed to maintain delta-neutral risk exposure for Uniswap V3 positions. However, based on our European LP option pricing model, we find that delta risk within the $(L,H)$ range differs from the delta derived from payoff-based calculations.
Taking the comparison between the orange and blue lines in Figure 4 as an example (since it is uncommon to persist with Uniswap V3 positions under unfavorable market conditions), we observe that when the price is at 0.9, the delta value of the orange line exceeds that of the blue line. This indicates that the actual delta risk at this point is higher than the delta risk derived from payoff-based pricing. Consequently, more delta needs to be hedged to achieve delta neutrality. Conversely, when the price is at 1.1, the actual delta risk associated with holding the Uniswap V3 position is lower. This suggests that hedging excessive delta based on payoff-based calculations is unnecessary.
@@ IV-A Delta Analysis
IV-B Gamma Analysis
By conducting a gamma analysis of the European LP pricing model versus the payoff-based pricing model, we can observe from Figure 5 that holding a Uniswap V3 position inherently involves gamma risk. Gamma risk refers to the situation where large gamma values cause significant fluctuations in delta with small changes in the underlying price $S$. In such scenarios, the delta hedging strategy becomes less effective, as even minor changes in $S$ can lead to substantial delta risk, necessitating more frequent adjustments to maintain delta neutrality. Conversely, when gamma values are small, delta remains relatively stable, reducing the need for frequent hedging.
As shown in the Figure 5, there is a significant difference between the gamma risk values derived from payoff-based modeling and those from the European LP pricing model. This highlights the higher requirements for managing gamma risk in Uniswap V3 positions, emphasizing the need for more refined risk control strategies.
IV-C Vega Analysis
@@
IV-D Greeks Comparison between different range order
After analyzing the unique patterns exhibited by the Greeks under Uniswap V3 positions, we aim to compare the Greeks of Uniswap V3 positions established with different $(L,H)$ ranges to uncover the relationship between Uniswap V3 pricing and the choice of $(L,H)$. The results are illustrated in the figure.
We set up three groups of Uniswap V3 range orders with price ranges of $(0.8,1.1)$, $(0.8,1.2)$, and $(0.8,1.3)$. The assumed market conditions are $r=0.05$, $C=0.2$, $\mu=0$, and $\sigma=0.7$. From Figure 7 ,we observe that as the pricing range widens, the European LP pricing model’s valuation results tend to increase. This is because a wider pricing range implies that the asset is expected to remain within the pricing range for a longer period, thereby extending the expected time for fee collection.
@@ IV-D Greeks Comparison between different range order
An interesting observation arises from the delta comparison in Figure 8 . We find that as the pricing range widens, the delta values generally increase. This is because when the pricing range is very wide, the proportional impact of price fluctuations on the exchange between the underlying asset and currency diminishes relative to the total position value. Consequently, the position value remains strongly correlated with the underlying asset price, resulting in a higher delta. Conversely, when the pricing range is narrow, even small price movements can trigger significant exchange activity, weakening the correlation between position value and underlying asset price, thus reducing delta.
@@
Although wider pricing ranges lead to larger delta risks, the absolute value of gamma decreases, as shown in Figure 9. This indicates that delta hedging becomes easier for wider pricing ranges compared to narrower ones. With a broader pricing range, delta changes more smoothly as the spot price fluctuates, simplifying the delta hedging process.
@@
Finally, as demonstrated in Figure 10, even with changes to the pricing range, vega remains strictly negative. This confirms that the relationship between price and volatility consistently exhibits a negative correlation.
@@
To validate the effectiveness of our Greeks, we need to perform numerical tests. Specifically, based on the Taylor expansion method, the following equation holds:
$\Delta V\approx\text{Delta}\times\Delta S+\frac{1}{2}\text{Gamma}\times\Delta S^{2}+\text{Vega}\times\Delta\sigma$
We can use two sets of different parameters, $S$ and $\sigma$, to calculate the valuation change of our Uniswap V3 position before and after market fluctuations (based on the present value derived from our pricing model). If the Greeks we calculate, when integrated into the right-hand side of the equation, can estimate the pricing with an acceptable level of accuracy, it will sufficiently demonstrate the validity of our Greeks.
The parameters we set are as follows, for the market before the fluctuation:
$C=0.2,r=0.05,\mu=0,\sigma=0.7$
$S=1,H=1.1,L=0.8$
Taking ETH as an example, under general circumstances, we assume a 5 % decrease in price and a 10 % increase in volatility itself. Using our pricing model, the valuation of the Uniswap V3 position before and after the market fluctuation is calculated as follows:
The Current Pv is: 1.02110809997664861.0211080999766486 1.02110809997664861.0211080999766486
The new PV is: 0.99304363754849550.9930436375484955 0.99304363754849550.9930436375484955
The Position Changing is $-0.028064462428153192$
The Greeks obtained through calculation, along with the corresponding Taylor expansion results, are as follows:
The Current delta is: 0.197906601375752440.19790660137575244 0.197906601375752440.19790660137575244
The Current gamma is: $-11.101273578173476$
The Current vega is: $-0.15958993211606654$
Taylor Expansion estimation is: $-0.034943217289629125$
The Current residential is: $-0.006878754861475933$
Thus, under this level of market changes, the pricing estimation derived from the Taylor expansion exhibits an approximate error of 0.6% compared to the actual valuation of the Uniswap V3 position.
IV-E The Impact of Volatility
The relationship between the volatility of the underlying asset and the pricing dynamics of the Uniswap V3 position is integral to our sensitivity analysis. Intuitively, we understand that according to the European option valuation model, the present value (PV) pricing should exhibit a monotonically negative correlation with volatility. This stems from the Automatic Market Maker (AMM) mechanism of Uniswap V3, which fundamentally behaves as a short volatility strategy. As the volatility increases, we anticipate a decrease in the valuation, signifying an inverse correlation.
However, the observed orange line in Figure 11deviates from this predicted relationship due to the dual pricing model for transaction fees, which has separate methodologies for calculating both upper and lower bounds. The rationale behind the lower bound is driven by the expectation of time until a transaction fee is collected, conceptualized through the theory of stopping times, which is directly proportional to volatility. Conversely, the upper bound assumes continuous trading throughout the transactional process, operating under the premise that serial transactions yield continuous fees, justifying the positive valuation associated with low volatility events.
We selected points with relatively low volatility, such as $\sigma$=0.02, to compare the pricing models of the two transaction fee calculation methods in terms of their present value (PV), as shown in Figure 12. In the PV-spot graph, the pricing model based on the upper bound of transaction fees consistently yields a higher value than the pricing model based on the lower bound of transaction fees.
@@ IV-E The Impact of Volatility
V future work
V-1 Hedging Strategies
Our pricing model is equipped to offer more accurate calculations of Greek risk measures such as delta, gamma, and vega. These calculations are crucial for developing sophisticated hedging strategies. By leveraging these measures, we can devise fully hedged strategies that maximize profitability while minimizing risk. Such strategies will not only benefit LPs by enhancing their returns and reducing risk but also contribute to the overall stability and efficiency of the DeFi ecosystem.
V-2 Adaptation to Uniswap v4
With the introduction of Uniswap v4, various new decentralized exchange (DEX) paradigms will be implemented, such as dynamic fee structures [7], which make $C$ a function of volatility. This simplification of the model allows LPs to adjust their positions less frequently in response to market conditions, making participation more accessible and efficient. Beyond dynamic fees, many potential enhancements and new models may emerge with Uniswap v4, requiring further adaptations and improvements to our frameworks.
V-3 Implied volatility
Our model is not only applicable for pricing but also serves as a tool for estimating implied volatility. Similar to the Black-Scholes-Merton (BSM) framework, we can infer the views of LPs on volatility through their positions. By aggregating these inferred volatilities, we can construct a liquidity surface that reflects the weighted perspectives of LPs.

### v3

@@ III Sensitivity Analysis
Our Greeks analysis will use two market parameter sets: one with high fees and high volatility (”good market”), and another with low fees and low volatility (”bad market”). Using the European LP option pricing model, we compare valuation differences between good/bad market conditions and payoff models across price levels, with a range order interval of (0.8, 1.2).
As shown in Figure 3, the European LP model reveals a significant gap between Uniswap V3’s current payoff and its present value, as the latter includes discounted future fees. In low volatility and fee scenarios, not withdrawing the position leads to value depreciation, with future fees failing to offset losses (green line). Conversely, in high volatility and fee ”good market” conditions, the position’s actual value exceeds its payoff (orange line).
@@
III-A Delta Analysis
In Uniswap V3 trading, delta hedging is often used to manage delta-neutral risk. However, our European LP option pricing model reveals that delta risk within $(L,H)$ the range differs from payoff-based delta.
For example, in Figure 4, the orange line (model delta) exceeds the blue line (payoff-based delta) at a price of 0.9, indicating higher actual delta risk and requiring more hedging. Conversely, at 1.1, model delta risk is lower, making excessive hedging unnecessary.
@@ III-A Delta Analysis
III-B Gamma Analysis
Gamma analysis of the European LP pricing model versus the payoff-based model (Figure 5) shows that Uniswap V3 positions inherently carry gamma risk. Large gamma values cause delta to fluctuate significantly with small changes in $S$, reducing delta hedging effectiveness and requiring frequent adjustments. Conversely, small gamma values stabilize delta, minimizing hedging needs.
Figure 5 reveals a significant gap between gamma risks from payoff-based and European LP models, highlighting the need for more refined risk control in Uniswap V3 positions.
III-C Vega Analysis
@@
III-D Greeks Comparison between different range order
[…]
@@ III-D Greeks Comparison between different range order
From Figure 7, wider ranges increase European LP valuations, as assets stay longer within the range, extending fee collection time. Figure 8 shows delta rises with range width, as price fluctuations impact exchanges less relative to total position value, strengthening price correlation. Conversely, narrow ranges weaken this correlation, reducing delta.
Figure 9 indicates that while wider ranges increase delta risk, gamma decreases, making delta hedging easier due to smoother delta changes. Finally, Figure 10 confirms vega remains negative across ranges, highlighting the negative price-volatility relationship.
III-E The Impact of Volatility
The relationship between underlying asset volatility and Uniswap V3 pricing is key to our sensitivity analysis. Intuitively, the European option model suggests PV should negatively correlate with volatility, as Uniswap V3’s AMM acts as a short volatility strategy. Higher volatility thus lowers valuation.
However, Figure 11’s orange line deviates due to the dual pricing model for fees. The lower bound reflects fee collection time, proportional to volatility via stopping time theory. The upper bound assumes continuous trading, justifying positive valuation in low volatility scenarios.
Figure 12 compares PV for $\sigma=0.02$. The upper-bound fee model consistently yields higher PV than the lower-bound model.
@@ III-E The Impact of Volatility
IV Application of model
In this section, we apply model to calculate the implied volatility and try to generalize the model to more specific AMM case.
IV-A Implied Volatility
Building upon the foundation of pricing research, we can also deduce the implied volatility of corresponding Uni V3 positions by inversely solving the pricing model.We can aggregate the implied volatility views of all Uni V3 investors at a cross-section, weighting them to observe the historical performance of the weighted implied volatility views. Notably, when we compare the historical performance of implied volatility with LVR model, we uncover some fascinating patterns.
The volatility inferred by LVR is based on the assumption of no arbitrage. We possess an instantaneous LVR formula, and we also understand that the collection of fees is approximately a linear model over an extremely short period. According to the [14], under the no-arbitrage assumption, the instantaneous LVR should equate to the instantaneous fees collected.
$C\lambda=\frac{\lambda\sigma^{2}}{4}$
Based on this established equality, the sigma term within the instantaneous LVR formula can be inversely calculated by other parameters in the equation, thereby yielding an implied volatility formula based on LVR.
$\sigma=2\sqrt{C}$
We collected historical transaction data for the ETH-USDC-005 trading pair on the Ethereum blockchain from June to December 2024 and created a comparative chart of implied volatility, the figure 13. We can observe that the historical performance of the weighted IV largely aligns with the trend of the IV inferred by LVR. Additionally, the graph presents unweighted and non-aggregated IV samples, revealing that the IV deduced by LVR predominantly resides at the upper echelon of the IV samples. This phenomenon can be attributed to our model’s assumption that future returns are subject to discounting. Consequently, our model inherently presumes the advantage of early boundary attainment for profit realization. As a result, the IV derived from our model is comparatively underestimated relative to the IV inferred by LVR.
Figure 13: IV histogram Comparison
While the implied volatility expressions are notably similar across the three models, key differences lie in their assumptions and specifics:
  1. 1.
    Lambert’s model assumes theta approximates a Dirac delta as time approaches zero, deriving the final option price. LVR also relies on infinitesimal time. Our model uses stopping time of a stochastic process to determine expiration, providing expected maturity time.
  2. 2.
    LVR and Lambert assume zero risk-free rate, equating future earnings to present value. Our model incorporates the risk-free rate for more accurate present value estimation of Uni V3 positions.
  3. 3.
    LVR and Lambert can only give the instantaneous value of Uni V3 position, in hence they can’t capture greeks. On the contrary, our model can easily calculate greeks to measure the risk.
  4. 4.
    LVR and Lambert focus on liquidity pool pricing, ignoring range order differences. We argue range orders reflect investor views on price and volatility, making price range relevant to Uni V3 pricing
  5. 5.
    LVR and Lambert assume zero-drift geometric Brownian motion. Our risk-neutral model allows flexible drift assumptions.
  6. 6.
    Like LVR, our model excels in scalability, as demonstrated in the next subsection.
We summarize three models as following table, RNP represent our model’s name risk-neutral-pricing model.
RNP Lambert LVR
Maturity stopping time 0 flexible
risk free flexible 0 0
greeks support non-support non-support
AMM range relevant irrelevant irrelevant
price process GBM 0 drift GBM 0 drift GBM
scalability other AMM none other AMM
IV-B Extend to other AMM case
IV-B1 Example of application in Uniswap V2
In the case of Uniswap V2, as the liquidity bounds (LH) approach zero and infinity, the stopping time for the euro-situation tends toward infinity, rendering the model ineffective. However, the modeling approach for the amer-situation can still be applied to Uniswap V2. This is because, as the duration of the position extends, the discounted value of the collected fees has a diminishing impact on overall returns, making fee collection less attractive to investors. The price range at which investors choose to exit their Uniswap V2 positions influences the expected duration of the position. Therefore, based on the price range at which investors actively close their positions, our model can still price the Uniswap V2 position.
As the formula represent:
$L_{1}<x<L_{2}$
$d=\frac{\ln L_{2}}{\sigma},d^{\prime}=d-x$
$c=\frac{\ln L_{1}}{\sigma},c^{\prime}=x-c$
So we have the formula:
$V_{E}$ $=LP_{L_{2}}^{V2\ case}\exp({\mu^{\prime}d^{\prime}})\frac{\sinh[c^{\prime}\sqrt{\mu^{\prime 2}+2r}]}{\sinh[(d-c)\sqrt{\mu^{\prime 2}+2r}]}$
$+LP_{L_{1}}^{V2\ case}\exp({-\mu^{\prime}c^{\prime}})\frac{\sinh[(d^{\prime}\sqrt{\mu^{\prime 2}+2r}]}{\sinh[(d-c)\sqrt{\mu^{\prime 2}+2r}]}$
This formula can solve the Uniswap V2 pricing problem.
IV-B2 Application in Uniswap V4: Dynamic fee example
Uniswap V4 has a unique mechanism called dynamic fee. Unlike V3’s constant fee. Dynamic fee can be setting more flexible. One kind of dynamic fee can be adjusted in real-time based on various market conditions, like volatility. It means we can regard fee as a variable that is directly influenced by volatility. Our model can easily adapt dynamic fee mechanism, by adjusting the internal parameter C of the model into a function of volatility. We believe our model will also be highly adaptable to the pricing of Uniswap V4 positions for variety hooks.
V Conclusion
In this paper, we have constructed a risk-neutral pricing model based on the stopping time theorem to price Uniswap V3 positions. Our model capture the value of each investor’s range order is predicated on the discount of all potential earnings during the period until they first exit the position. From this hypothesis, we regard the Uniswap V3 position as an option pricing problem and have derived several analytical solution based on euro-situation and amer-situation, along with the corresponding Greeks based on these pricing formulas. The Greeks can be utilized to measure the risk exposure faced by investors holding this position. We have also employed MC simulations to verify the accuracy of our formula. This pricing model successfully delineates the relationship between the value of a Uniswap V3 position, its range setting, and the market’s volatility. Through this model, investors can discern the implied volatility perspective underlying their Uni V3 positions and compare it with the prevailing liquidity views in the market. Moreover, with an accurate estimation of risk exposure, investors can more readily manage their risks and steer clear of undesired exposures.